OpenStudy (anonymous):

Can someone wlak me through this...Express sinx +sqrt3 cosx in terms of Asin(x+c)

myininaya (myininaya):

so we have \[1 \cdot \sin(x)+\sqrt{3} \cdot \cos(x)\] recall \[\cos(C)\sin(x)+\sin(C)\cos(x)=\sin(x+C)\] and if we multiply both sides of that by A we have \[A \cos(C)\sin(x)+A \sin(C)\cos(x)=A \sin(x+C)\] so looking at our expression and comparing to the expression in this drawing box:  compare those in that box what do we need what to be?